OpenStudy (anonymous):

For f(x) = 0.01(2)x, find the average rate of change from x = 3 to x = 8.

OpenStudy (anonymous):

@jojo12g

OpenStudy (anonymous):

@King.Void.

OpenStudy (anonymous):

@MrCoolGuy

Vocaloid (vocaloid):

rate of change = [f(8)-f(5)]/(8-5)

MsBrains (ms-brains):

Can you please snip the whole problem, and paste the link. So I can see the original and help better

OpenStudy (jojo12g):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (anonymous):

@Preetha

OpenStudy (jojo12g):

A because 0.01(2)4=0.08 @madison_taylor914

MsBrains (ms-brains):

I'm not sure about this one... Sorry Madison.

OpenStudy (anonymous):

Which of the following options results in a graph that shows exponential decay? f(x) = 0.6(2)^x f(x) = 3(0.7)^x f(x) = 0.4(1.6)^x f(x) = 20(3)^x

OpenStudy (anonymous):

@Ms-Brains

OpenStudy (jojo12g):

A because 0.01(2)4=0.08 @madison_taylor914

OpenStudy (anonymous):

thats not what its asking @jojo12g

OpenStudy (mrcoolguy):

im not sure on this one

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[f(x)=0.1\times 2^x\] right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

you need to compute \[f(8),f(3)\] and then take the slope \[\frac{f(8)-f(3)}{8-3}\]

OpenStudy (misty1212):

the function is \[f(x)=0.01\times 2^x\] so \[f(8)=2.56\]

OpenStudy (misty1212):

\[f(3)=0.01\times 2^3=0.01\times 8=0.08\]

OpenStudy (anonymous):

so 2.48 final answer?

OpenStudy (misty1212):

no you still have to divide by 5

OpenStudy (anonymous):

.496

OpenStudy (lochana):

yes. that's the answer

OpenStudy (anonymous):

thanks guys

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

can you help me with one more? @misty1212